OpenStudy (anonymous):

How can I differentiate the function y=6/ln (x) ?

OpenStudy (anonymous):

Use the quotient rule: Your numerator is f(x), the denominator is g(x). So f(x)=6, g(x)=ln(x). The Quotient Rule is: ( (g(x) * f'(x)) - ((f(x) * g'(x)) )/ g(x)^2 f(x) = 6 => f'(x)=0 g(x) = lnx => g'(x)=1/x Use the formula to find your answer.